The cost for a long-distance telephone call is $0.36 for the first minute and $0.21 for each additional minute or portion thereof. Write an inequality representing the number of minutes a person could talk without exceeding $3.
Answer by stanbon(75887)

The cost for a long-distance telephone call is $0.36 for the first minute and $0.21 for each additional minute or portion thereof. Write an inequality representing the number of minutes a person could talk without exceeding $3.
-----------------
Cost(x minutes) = 0.36 + 0.21(x-1)
INEQUALITY:
0.36 + 0.21x - 0.21 <= 3
0.15 + 0.21x <= 3
0.21x <= 2.85
x <= 13.57minutes
===============
Cheers,
Stan H.